OpenStudy (wesdg1978):

How do you: Evaluate the lim x-->pi/2 e^tant

OpenStudy (anonymous):

Check what you wrote, is it suppose to be e^tan(x)?

OpenStudy (noelgreco):

What does the tangent do at pi/2?

OpenStudy (wesdg1978):

Maybe @mathis1, the paper I have has a t not x

OpenStudy (anonymous):

If that is the case, there is no x to plug in, and the limit x->pi/2 = e^tant

OpenStudy (johnweldon1993):

If it really just has a t....then the function isn't affected no matter where x approaches...so it would just be e^tant

OpenStudy (anonymous):

so it is lim t->pi/2

OpenStudy (wesdg1978):

I copied it correctly, it is lim x-->pi/2 e^tant

OpenStudy (wesdg1978):

If there is a typo, it's on the worksheet I copied from.

OpenStudy (anonymous):

Well if the limit and the tan(x) are not the same variable, it is like I said before, the variables are not the same, so the limit is just the e^tant. If they are the same variable, then you need to think, what part of the equation is the most greatly affected, in this case the tanx. So, what does tanx do at pi/2?

OpenStudy (noelgreco):

\[\lim _{t \rightarrow \frac{ \pi }{ 2 }}e ^{\tan t}\] Is that it?

OpenStudy (anonymous):

A two sided limit does not exist... \(\tan t\to\infty\) as \(t\to\pi/2^-\) so approaching from the left should give us \(e^{\tan t}\to\infty\). From the right, however, \(\tan t\to-\infty\) and thus \(e^{\tan t}\to 0\)

OpenStudy (wesdg1978):

Honestly, I'm not sure what tan x does at pi/2.

OpenStudy (wesdg1978):

@NoelGreco I'm not sure if the paper I have has a typo or if it is supposed to be x and t

OpenStudy (noelgreco):

It's a typo because t could not care less what x is doing.

OpenStudy (anonymous):

@wesdg1978 \(\cos t\to0\) as \(t\to\pi/2\) yet \(\sin t\to1\) so \(\tan t=\sin t/\cos t\to\pm\infty\) depending on whether you approach it from where \(\cos t<0\) or \(\cos t>0\)

OpenStudy (noelgreco):

\[\tan x=\frac{ \sin x }{ \cos x }\] What does the denominator approach as the variable goes to zero.

OpenStudy (wesdg1978):

1

OpenStudy (noelgreco):

Sorry. What does the denominator approach as the variable goes to pi/2?

OpenStudy (wesdg1978):

0

OpenStudy (noelgreco):

Good.

OpenStudy (noelgreco):

Tan pi/2 then is what?

OpenStudy (wesdg1978):

If the denominator is 0, then it would be undefined, right?

OpenStudy (noelgreco):

Yes, but for purposes of finding a limit, let's say infinity. Now we have \[e^{\infty}\]

OpenStudy (noelgreco):

What's the limit?

OpenStudy (noelgreco):

I have to turn in for the night Any last thoughts?

OpenStudy (wesdg1978):

My internet went out for like 5 minutes. Is it e^infinity is infinity

OpenStudy (wesdg1978):

My apologies, my computer locked up and my internet went out.

OpenStudy (noelgreco):

Bingo! Best answer is DNE (does not exist) the limit is +inf from the left, and -inf from the right. Learn those trig graphs! It'll pay off.

OpenStudy (wesdg1978):

Thank you very much!

OpenStudy (noelgreco):

Limits, trig, and a curious student? - my pleasure.

OpenStudy (wesdg1978):

After this test tomorrow, I'm going to go back and refresh memory with trig, any suggestions on the best way to learn trig graphs @NoelGreco ?